In a titration experiment, 9.3 mL of an aqueous H2SO4 solution was titrated with 0.3 M NaOH solution. The equivalence point in the titration was reached when 9.9 mL of the NaOH solution was added. What is the molarity of the H2SO4 solution?
Answer #1

Balanced chemical equation is:

2 NaOH + H2SO4 ---> Na2SO4 + 2 H2O

Here:

M(NaOH)=0.3 M

V(NaOH)=9.9 mL

V(H2SO4)=9.3 mL

According to balanced reaction:

1*number of mol of NaOH =2*number of mol of H2SO4

1*M(NaOH)*V(NaOH) =2*M(H2SO4)*V(H2SO4)

1*0.3*9.9 = 2*M(H2SO4)*9.3

M(H2SO4) = 0.1597 M

Answer: 0.16 M